OpenStudy (anonymous):

Evaluate x/y, for x =2/3 and y = 5/6 Please help asap

OpenStudy (anonymous):

X/Y= 2/3/5/6 =2/3*6/5 =4/5

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Is (5, 17) a solution of the equation y = 4x + 3? /why?

OpenStudy (anonymous):

No. Because if you plug in x=5 into the equation, you dont get y= 17

OpenStudy (anonymous):

Thanks ..sorry im not good with algebra

OpenStudy (anonymous):

Not a problem.

OpenStudy (anonymous):

can i ask you another?

OpenStudy (anonymous):

Yeah, last one for now though cos i gotta go

OpenStudy (anonymous):

Suppose U = {1, 2, 3, 4, 5, 6, 7, 8} is the universal set and P = {1, 2, 3, 4}. What is P'?

OpenStudy (anonymous):

and okay

OpenStudy (anonymous):

P' means P complement. This means things in P that are not in U ={5,6,7,8}

OpenStudy (anonymous):

If you have anymore problems, just tag me, ill help later

OpenStudy (anonymous):

Ok thank you so much!